Evaluate โˆซf ๐‘œ๐‘ฃ๐‘’๐‘Ÿ ๐‘ where ๐‘“( ๐‘ง )= ๐‘ฅ^2 + ๐‘–๐‘ฆ^2 where c is given by ๐‘ง (๐‘ก )= ๐‘ก^2 + ๐‘–๐‘ก^2, 0 โ‰ค ๐‘ก โ‰ค 1

1
Expert's answer
2022-03-20T06:43:38-0400

"\\int_C{f}=\\left[ \\begin{array}{c}\tz\\left( t \\right) =t^2+it^2\\\\\tdz\\left( t \\right) =\\left( 2t+2it \\right) dt=2\\left( 1+i \\right) tdt\\\\\tf\\left( z \\right) =x^2+iy^2=\\\\\t=\\left( t^2 \\right) ^2+i\\left( t^2 \\right) ^2=\\\\\t=\\left( 1+i \\right) t^4\\\\\\end{array} \\right] =\\int_0^1{\\left( 1+i \\right) t^4\\cdot 2\\left( 1+i \\right) tdt}=\\\\=2\\left( 1+i \\right) ^2\\int_0^1{t^5dt}=\\frac{2\\cdot 2i}{6}=\\frac{2}{3}i"